### v4

# Identification of Craving Maps among Marijuana Users via the Analysis of Functional Brain Networks with High-Order Attention Graph Neural Networks

Abstract
The excessive consumption of marijuana can induce substantial psychological and social consequences. In this investigation, we propose an elucidative framework termed high-order graph attention neural networks (HOGANN) for the classification of Marijuana addiction, coupled with an analysis of localized brain network communities exhibiting abnormal activities among chronic marijuana users. HOGANN integrates dynamic intrinsic functional brain networks, estimated from resting-state functional magnetic resonance imaging (rs-fMRI), using long short-term memory (LSTM) to capture temporal network dynamics. We employ a high-order attention module for information fusion and message passing among neighboring nodes, enhancing the network community analysis. Our model is validated across two distinct data cohorts, yielding substantially higher classification accuracy than benchmark algorithms. Furthermore, we discern the most pertinent subnetworks and cognitive regions affected by persistent marijuana consumption, indicating adverse effects on functional brain networks, particularly within the dorsal attention and frontoparietal networks. Intriguingly, our model demonstrates superior performance in cohorts exhibiting prolonged dependence, implying that prolonged marijuana usage induces more pronounced alterations in brain networks. The model proficiently identifies craving brain maps, thereby delineating critical brain regions for analysis.
Keywords Marijuana  $\cdot$ fMRI  $\cdot$ Graph neural network (GNN)  $\cdot$ Multigraph classification  $\cdot$ Addiction prediction
1 Introduction
The issue of substance use disorder (SUD) has become a pressing concern in the United States, with the legalization of non-medical marijuana usage. Marijuana-related disorder accounts for a substantial proportion of individuals seeking treatment for drug use disorders due to the high global prevalence of marijuana use [1]. Heavy marijuana consumption has been found to impact brain function and cognition  [2, 3, 4]. In the traditional view, addiction is associated with brain abnormalities in the nucleus accumbens, prefrontal cortex, and amygdala, extensively investigated through neuroimaging techniques [5, 6, 7]. The regional structural or functional changes among patients with SUD may influence addictive behavior by altering reward processing, decision-making, and emotional regulation [8]. However, current research indicates that addiction is not merely a localized brain abnormality but a disruption at the network level. Interacting neural assemblies or brain regions constitute brain networks, and disruptions in connectivity between different regions may crucially impact addictive behavior. Addiction is thus viewed as a network-wide phenomenon with altered functional connectivity patterns, as indicated by network analysis based on fMRI [9, 10, 11, 12]. Recent studies have shown that leveraging fMRI-derived brain network analysis can offer new insights into the relationships between underlying brain connectivity alterations and the manifested characteristics of addictive behavior [13, 14, 15]. These characteristics include the age of onset of marijuana use, environmental influences, symptoms of addiction [14], and predictions of marijuana withdrawal and personalized treatment [13].
Researchers have developed several interpretable predictive frameworks based on classical machine learning models to identify individuals at risk of SUD. These frameworks, which utilize algorithms like support vector machines (SVMs) and random forests (RFs), can predict the likelihood of SUD based on different age groups [16]. The interpretable machine learning frameworks identified risk factors for SUD by investigating gender differences, personality characteristics, and cognitive abilities [17, 18, 19].
Constructing brain connectomes using functional or structural neuroimaging modalities has become one of the most pervasive frameworks for brain imaging analysis [20]. In recent years, graph neural networks (GNNs) have been employed to analyze brain networks for neurological disorder classification and subtyping [20]. These functional brain networks are constructed from correlation maps of fMRI, where each vertex represents a region of interest (ROI), and edges represent the connection strength between different ROIs  [21, 22, 23, 24]. Researchers have utilized a multimodal approach of diffusion tensor imaging (DTI) and fMRI to diagnose mental network disorders [22], and have also built functional brain networks to classify autism spectrum disorder (ASD)  [25, 26, 27]. Moreover, Schizophrenia can be identified with salient regions based on its topological structure using SVM and Graph Convolutional Network (GCN), as demonstrated by six rs-fMRI datasets. [28].
However, fewer studies have conducted functional brain network analysis based on GNN for long-term marijuana users to (1) improve the classification accuracy and (2) delineate the related craving maps or explain alterations of brain networks. To bridge the research gap, in this study, we propose an interpretable high-order attention graph attention neural networks (HOGANN) model for analyzing the abnormal brain activities demonstrated in long-term marijuana users compared to healthy controls (HC). The main contributions of this study can be summarized as follows: (1) proposing to utilize graph deep learning methods to classify long-term marijuana users, (2) leveraging the sequential multigraphs in the fMRI time series of marijuana users, (3) integrating the functionality of high-order nodes for information fusion and message passing into our model, and (4) providing improved clustering analysis for identifying craving maps among long-term marijuana users. Furthermore, we compared multiple benchmark methods and evaluated the effectiveness of the proposed HOGANN. By training the model and reconstructing weighted brain networks, we found that the HOGANN is more effective at analyzing differences between long-term marijuana users and non-users as characterized by functional brain connectivity networks.
2 Method
2.1 Problem Definition
We define the classification of SUD for each subject as a supervised multigraph classification problem, where each instance graph is constructed from a sliding window of the resting state fMRI (rs-fMRI) time series. Specifically, we denote the multivariate time-series data for all individuals as $\{\textbf{X}^{(i)}\}_{i=1}^{N}$ with each individual’s data being $\textbf{X}^{(i)}\in\mathbb{R}^{v\times T}$, where $v$ represents the number of ROIs, and $T$ is the time series length. We segment these non-overlapping time series into multiple segments, each with a length of $T^{\prime}$. This results in a total of $\left|K\right|$ segments for each subject. Subsequently, we construct the multigraph $\{\textbf{G}^{(i)}(E,V;t)\}_{t=1}^{K}$ = $\{\textbf{A}^{(i)}(t)\}_{t=1}^{K}$ from the segmented subsequences, where the functional connectivity is based on the k-nearest neighbor (k-NN) graph of each ROI for each subject, where $E$ and $V$ denotes edges and vertex (region) set, respectively. Each adjacency matrix $\textbf{A}^{(i)}(t)\in\mathbb{R}^{v\times v}$ corresponds to the $k$th segment in the $t$th time window denoted $X^{(i)}(t),(i\in 1,\cdots,N,t\in 1,\cdots,K)$, and corresponding class labels $Y^{(i)}$ for the multigraph classification task.
2.2 Model Architecture
The overall framework of HOGANN, as shown in Fig. 1encompasses a multigraph classification architecture with two sub-models: (1) the high-order graph attention neighbor mixing time series model, which captures features from multigraphs of the same subject, and (2) the sequential dynamic graph learning model using LSTM to learn the temporal interdependence of graphs. The high-order graph attention neighbor mixing module employs a higher-order attention mechanism with a neighborhood mixing model to learn the relationship between ROIs after segmenting dynamic time series in the same subject. Meanwhile, the proposed LSTM sub-model can better extract local features from the sequential graphs to improve distinguishability between marijuana subjects and normal controls. The overall framework enables concatenating features from the latent spaces of the two sub-models leading to improved model performance.
Figure 1: The overall framework of the HOGANN utilizes two sub-models to perform fusion learning on fMRI time series. The high-order graph attention neighbor mixing model combines hopping to conduct high-order attention for message passing. Sequential dynamic graph learning uses LSTM to learn the temporal sequence graph and discern differences between instances.
2.2.1 High-order graph attention neighbor mixing
To learn dynamic topological brain networks from multigraphs, we utilize one-hop connections for local graph node message passing with a MixHop layer [29]. The formulated equation of the classical $l$-th graph hidden convolutional layer can be defined as follows:
$\textbf{M}^{(l+1)}=\sigma(\hat{\textbf{A}}\textbf{M}^{(l)}\textbf{W}^{(l)}),$ (1)
where normalized adjacency matrix is given by $\hat{\textbf{A}}=\textbf{D}^{-\frac{1}{2}}(\textbf{A}+\textbf{I}_{N})\textbf{D}^{-\frac{1}{2}}$, and $\textbf{A}+\textbf{I}_{N}$ denotes the self-loop connection; $D_{ii}=\sum_{i}A_{ii}$ presents the degree matrix of A, W is a learnable weight matrix in the non-linear activation function $\sigma(\cdot)$. By incorporating the neighborhoods of high-order nodes, the MixHop layer facilitates higher-order latent feature learning enables learning from their neighbors by incorporating the $j$-th power of the self-adjacency matrix $\hat{\textbf{A}}^{j}$, and the two-hop delta operator learns for various neighborhood distances by subtracting different node features. The MixHop graph convolutional layer can be defined as
$\textbf{M}^{(l+1)}=\mathop{\Big{\|}}\limits_{j\in P}\sigma(\hat{\textbf{A}}^{j}\textbf{M}^{(i)}\textbf{W}_{j}^{(i)}),$ (2)
[…]
Attention with MixHop fusion: To effectively integrate temporal graph features and the strength of connectivity paths between nodes through the attention mechanism, we utilized both the MixHop architecture employing high-order message passing and the GAT module for fusing multigraph node features, as depicted in Fig. 2. The GAT layer, with graph attention mechanisms, aggregate neighborhood node features with weighted attention score, while the MixHop layer enhances the connectivity path from source node to high neighbor node of functional connectivity [30]. (Note: first-degree to $k$-degree proximity). We denote the node embedding $\textbf{H}^{0}=\{\vec{h}^{0}_{1},\vec{h}^{0}_{2},...,\vec{h}^{0}_{n}\}$ from the first graph attention layer and learn the neighbors $\mathcal{N}_{i}$ of ROI nodes $v_{j}$. We can update the parameterized feature vectors $h_{i}$ and the learnable weights $\textbf{W}^{0}\in\mathbb{R}^{n\times d}$ using the LeakyReLU function to further calculate the importance attention score as Eq. (3)
$\alpha_{ij}=\frac{exp\left(LeakyReLU(\vec{a}^{T}_{ij}\left[\textbf{W}\vec{h}_{i}\mathbin{\|}\textbf{W}\vec{h}_{j}\right]\right)}{\sum_{k\in\mathcal{N}_{i}}exp\left(LeakyReLU(\vec{a}^{T}_{ij}\left[\textbf{W}\vec{h}_{i}\mathbin{\|}\textbf{W}\vec{h}_{k}\right]\right)},$ (3)
The attention mechanism generates an importance attention score $\alpha_{ij}\in\mathbb{R}$ for each edge between ROIs, enabling the neighboring vertex to better focus on the corresponding node at the edge. Subsequently, we can calculate the features representation $\hat{h}$ of the corresponding nodes through the attention coefficients after a non-linear transformation, as follows:
$\hat{h}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\textbf{W}\vec{h}_{j}\right),$ (4)
where $\sigma(\cdot)$ denotes a nonlinear transformation function. In the second layer of the sub-model, we enhance the attention mechanism and the capability for high-order feature fusion within the overall network, as illustrated in Fig. 2. We concatenate embedding vectors $m$ and $\hat{h}$ from the MixHop and GAT layers as the input to the second-layer graph attention layer. However, traditional attention weights for each edge are primarily effective for learning first-order node connectivity, whereas attention MixHop fusion excels at learning second-order neighbor proximity pathways. The concatenated embedding is utilized as the aggregation input for the second-layer $\textbf{GAT}_{l+1}$, as follows:
$\textbf{H}^{a}=\textbf{GAT}_{l+1}(\phi(\left[\textbf{M}\oplus\hat{\textbf{H}}\right])),$ (5)
where M and $\textbf{H}^{a}=\left[\hat{h}_{1}^{a},\hat{h}_{2}^{a},...,\hat{h}_{k}^{a}\right]$ are output features vectors, $\oplus$ present concatenation with corresponding output channel dimensions, $\phi(\cdot)$ denotes graph global mean pooling. The $\frac{1}{\left|\mathcal{V}\right|}\sum_{i\in\mathcal{V}}\hat{h}^{a(l)}_{i}$ indicates averaging node features across the node dimension $\left|\mathcal{V}\right|$.
2.2.2 Sequential graph learning
Figure 2: The proposed first sub-model contains MixHop and a graph attention layer, wherein the second layer of message passing, concatenating different powers of $\hat{A}$, can enhance the depth of attention in the neighborhood node connectivity in the pathway.
The fMRI signals of the brain’s ROI exhibit variations in dynamic network representation over time $T$ for each instance. The second sub-model was designed to extract dynamic features of functional graphs with cross-sequential dependencies using a batch of $b$ multigraphs ${\{\textbf{A}^{(i)}}\}_{i=1}^{N}$, where $b\subseteq|K|\times|N|$. LSTM networks have found extensive application in handling sequential data, particularly in scenarios where extracting interrelated sequential time series at various time points is necessary. Starting from time step t, we can formulate the sequence learning representation, which comprises the input gate $t$, we can formulate the sequence learning representation, which comprises the input gate $\textbf{I}_{t}$, forget gate $\textbf{F}_{t}$, output gate $\textbf{O}_{t}$, cell gate $\textbf{C}_{t}$, and output gate $\textbf{H}^{g}$, as depicted:
Figure 3: The proposed second sub-model, LSTM, aims to train the graph of each batch instance and learn representations of spatial-temporal dependencies for different subjects.
$\textbf{I}_{t}$ $=\sigma(\textbf{W}_{i}\cdot\textbf{GATconv}(\textbf{A}^{(i)}(t))+\textbf{W}_{i}\cdot\textbf{GATconv}(H_{t-1}))+b_{i}$ (6)
$\textbf{F}_{t}$ $=\sigma(\textbf{W}_{f}\cdot\textbf{GATconv}(\textbf{A}^{(i)}(t))+\textbf{W}_{f}\cdot\textbf{GATconv}(H_{t-1}))+b_{f}$ (7)
$\textbf{O}_{t}$ $=\sigma(\textbf{W}_{o}\cdot\textbf{GATconv}(\textbf{A}^{(i)}(t))+\textbf{W}_{o}\cdot\textbf{GATconv}(H_{t-1}))+b_{o}$ (8)
$\tilde{\textbf{C}}_{t}$ $=tanh(\textbf{W}\cdot\textbf{GATconv}(\textbf{A}^{(i)}(t))+\textbf{W}\cdot\textbf{GATconv}(H_{t-1}))$ (9)
$\textbf{C}_{t}$ $=tanh(\textbf{I}_{t}\cdot\tilde{\textbf{C}}_{t}+\textbf{F}_{t}\cdot\textbf{C}_{t-1})$ (10)
H $=\textbf{O}_{t}\cdot tanh(\textbf{C}_{t})$ (11)
where $\textbf{W}{i}$, $\textbf{W}_{f}$, $\textbf{W}{o}$ and $\textbf{W}_{c}$ represent learnable weights at different gates and bias vectors $b_{i}$, $b_{f}$, and $b_{o}$. By taking the output from the LSTM layer, we obtain the final instance graph representation:
$h^{\prime}_{1},h^{\prime}_{2},...,h^{\prime}_{k}=\textbf{LSTM}(h_{1},h_{2},...,h_{k}),$ (12)
where $h^{\prime}_{1},h^{\prime}_{2},...,h^{\prime}_{k}$ are latent vectors in $t$-th time step from LSTM modeling as illustrated in Fig. 3. Then, we can combine temporal graph features and use graph mean pooling $\phi(\cdot)$ to generate instance representation $\textbf{H}^{\prime}=\phi(\left[h_{1}^{\prime}\lVert h_{2}^{\prime}\lVert h_{3}^{\prime}...\lVert h_{k}^{\prime}\right])$. To better aggregate two sub-model embeddings, we combine two representations of $\textbf{H}^{a}$ and $\textbf{H}^{\prime}$ for the final embedding output as shown:
$\textbf{H}^{c}_{i}=\textbf{H}^{a}+\textbf{H}^{\prime}.$ (13)
where $\textbf{H}^{c}_{i}\in\mathbb{R}^{B\times C}$ represents the final output, and C denotes the final prediction class. This prediction class is a combination of outputs from the two sub-models and is utilized to integrate information contributed by both perspectives provided by the sub-models.
2.2.3 Brain network reconstruction and community detection
Functional connectivity assesses brain connections between various regions. To enhance network connectivity and community structure, we utilize the final output embedding, denoted by $\textbf{H}^{c}$, for network reconstruction. This is accomplished through an inner product operation, as depicted in Eq. (14):
$\textbf{B}=\sigma(\textbf{H}^{c}\cdot\textbf{H}^{c^{T}}),$ (14)
where $\textbf{B}\in\mathbb{R}^{n\times n}$ represents the reconstructed matrix for ROIs. We utilize $\sigma(\cdot)$, for a nonlinear transformation to construct a weighted reconstructed matrix. Subsequently, for each subject’s connectivity matrix, denoted by $\textbf{A}^{(i)}_{r}(t)$ (where $i$ represents the subject and $r$ denotes the loop iteration), we perform element-wise multiplication with the weighted reconstructed matrix, $\textbf{B}_{r}$. This computation is shown in Eq. (15) and results in the final average weighted functional connectivity matrix.
$\textbf{W}^{FC}=\frac{1}{N}\sum_{t=1}^{N}\textbf{A}^{(i)}_{r}(t)\odot B_{t},$ (15)
After obtaining the average weighted functional connectivity $\textbf{W}^{FC}$, we further analyze the brain community structure for community detection, which $\odot$ denotes the element-wise product. The concept of community detection in graphs is based on the analysis of edge density. The Girvan-Newman (GN) modularity algorithm is utilized to ascertain whether based on the idea that there is a higher degree of similarity between nodes within a community and a lower degree of similarity between nodes outside the community [31, 32]. The GN algorithm follows these steps: (1) calculate the edge betweenness for each edge; (2) remove the edge with the highest edge betweenness; (3) recalculate the edge betweenness for the remaining edges in the network; (4) repeat steps (3) and (4) until any vertex in the network becomes a community. We can compute modularity $M$ and define community $C$ in the graph as:
$M=\frac{1}{2m}\sum_{ij}(W^{FC}_{ij}-P_{ij})\delta(C_{i},C_{j}).$ (16)
where $m$ represents the total number of edges in the graph, and $B$ denotes the adjacency matrix. We compute the expected number of edges $P_{ij}=p=2m/\left[n(n-1)\right]$ between $i$ and $j$ in Eq. (16). The $\delta$-function, considering vertices $i$ and $j$ belonging to the same community ($C_{i}=C_{j}$), aids in clustering within the community structure and assessing modularity.
2.2.4 Objective function
We can employ the output embedding in Eq. (5) and Eq. (13) throughout the training phase. Eq. (17) and Eq. (18) to execute a binary classification task, incorporating the final non-linear transformation $\sigma(\cdot)$ as follows:
$\hat{\textbf{Y}}_{c}=\sigma(\textbf{W}_{c}^{T}\textbf{H}_{i}^{c}+b_{c}),$ (17)
$\hat{\textbf{Y}}_{MixGAT}=\sigma(\textbf{W}_{a}^{T}\textbf{H}_{i}^{a}+b_{a}),$ (18)
where $\textbf{W}{c}$ and $\textbf{W}{a}$ represent the trained parameters, and $b_{c}$ and $b_{a}$ denotes biases. Subsequently, our objective is to minimize the overall loss function derived from the losses of two sub-models. The first sub-model pertains to the MixHop attention loss, denoted as $\mathcal{L}_{MixGAT}$, while the second sub-model involves the combination fusion model loss from Eq. (19) output embedding, denoted as $\mathcal{L}_{c}$. The combined loss can be expressed as:
$\mathcal{L}=\mathcal{L}_{MixGAT}(\textbf{Y},\hat{\textbf{Y}}_{MixGAT})+\mathcal{L}_{c}(\textbf{Y},\hat{\textbf{Y}}_{c}).$ (19)
where Y represents the ground truth label, while $\hat{\textbf{Y}}_{MixGAT}$ and $\hat{\textbf{Y}}_{c}$ represent the class predictions for marijuana users and healthy controls, respectively. By combining two $\mathcal{L}_{MixGAT}$ and $\mathcal{L}_{c}$, we can better leverage the fusion information in the total loss $\mathcal{L}$ not just sub-model loss.
3 Results
3.1 Dataset
  • Marijuana-323: In this study, we extended the prior work of Kaustubh R. et al  [33]. The research involved the collection of data from two datasets of magnetic resonance imaging (MRI) images comprising 125 and 198 participants, recruited either from the non-medical marijuana-seeking community or from hospitalized patients. Before and after the MRI scans, participants completed the marijuana craving questionnaire [34], marijuana withdrawal checklist, and marijuana problem investigation  [35]. The data underwent reprocessing from 3T imaging fMRI data, and an average time series of 90 ROIs was calculated for each subject based on the Stanford altas  [36]. In our final training and testing, there were 195 long-term marijuana users and 128 healthy control individuals.
  • HCP dataset: To assess the effectiveness of HOGANN, we evaluated an external dataset sourced from the human connectome project (HCP) S1200 [37], comprising 1096 rs-fMRI scans of young adults. The S1200 constitutes the major releases from 2012 to 2015 from Washington University in St. Louis, Missouri, United States. All subjects were recruited and approved by the Institutional Review Board (IRB) review committee at Washington University in St. Louis. Detailed information regarding HCP data preprocessing can be accessed from the project’s website(https://www.humanconnectome.org/). The HCP dataset encompasses 598 individuals who have used marijuana, with an average age of 28.76 ± 3.69, and 493 individuals who have never used marijuana, with an average age of 28.82 ± 3.72. This data preparation utilized each subject’s first session and excluded 5 rs-fMRIs containing fewer than 1200 frames. The cortical surfaces of each subject were partitioned into 22 brain regions serving as ROIs  [38], and blood oxygenation level-dependent (BOLD) time series signals were normalized using z-scores  [19].
3.2 Baseline methods
We conducted comparisons among linear methods, such as logistic regression (LR) and SVM [39], and graph-based methods with some state-of-the-art GNNs, and graph-based methods, including GNNs, including GCN, GraphSAGE [40], GAT, GraphSAGE, and BrainGNN [21] based on brain network construction pipelines. Linear classifiers were evaluated using both L1 and L2 regularization terms consistent with the original paper.
  • LR with penalty is a linear classification method that uses Lasso regularization to prevent overfitting. The penalty in L1 or L2 regularization is determined by the sum of the absolute values or the sum of the absolute values or the sum of squares of the feature weights. This technique drives the weights of specific features toward zero, thereby promoting feature selection.
  • SVM with penalty is a supervised learning algorithm that will be used for classification and regression analysis tasks. The combination of L1 and L2 penalty terms is referred to as Elastic Net. This is achieved by adding a linear combination of L1 and L2 penalty terms and adjusting the corresponding mixing parameter.
  • GCN is a semi-supervised graph convolutional network model designed to learn node representations by leveraging neighboring information. It is commonly applied to learning graph structural data.
  • GAT extends GNNs by introducing an attention mechanism that enables dynamic adjustment of node relationship weights during the learning process.
  • GraphSAGE is an inductive graph convolutional network capable of predicting unseen nodes. By employing a sampling strategy, the model learns node embeddings based on neighboring nodes’ features rather than relying on the entire graph.
  • BrainGB  [20] BrainGB is a platform founded on GNN methods. It facilitates predictions across functional and structural neuroimaging modalities. GNNs are employed to model relationships between different brain regions
Table 1: Overall classification performance
Dataset Model AUC (%) Accuracy (%) Precision (%) Recall (%)
L1 LR $77.7$ $77.2$ $78.0$ $77.7$
L2 LR $75.4$ $72.8$ $75.2$ $75.4$
L1 SVM $79.2$ $77.9$ $79.1$ $79.2$
L2 SVM $77.6$ $75.6$ $77.5$ $77.7$
Marijuana-323 GCN $77.7$ $71.5$ $71.9$ $71.5$
GAT $81.0$ $73.8$ $74.0$ $73.9$
GraphSAGE $77.3$ $74.6$ $75.4$ $74.6$
BrainGB $83.4$ $75.0$ $73.6$ $69.3$
HOGANN (w/o MixHop) $81.9$ $76.8$ $76.8$ $76.9$
HOGANN (w/o LSTM) $85.1$ $76.2$ $76.1$ $76.2$
HOGANN 85.1 80.7 81.0 80.7
L1 LR $62.8$ $62.7$ $63.3$ $62.7$
L2 LR $64.9$ $64.8$ $65.2$ $64.8$
L1 SVM $64.6$ $64.7$ $64.7$ $64.7$
L2 SVM $69.6$ $60.9$ $64.5$ $60.9$
HCP GCN $49.9$ $50.5$ $25.5$ $50.5$
GAT $65.6$ $68.2$ $68.3$ $68.2$
GraphSAGE $54.4$ $53.9$ $55.7$ $53.9$
BrainGB $54.4$ $54.5$ $52.4$ $51.9$
HOGANN (w/o MixHop) $70.4$ $64.8$ $66.0$ $64.8$
HOGANN (w/o LSTM) $66.3$ $67.2$ $67.3$ $67.2$
HOGANN 75.5 71.2 71.7 71.7
3.2.1 Linear classifier
We compared the proposed linear classifier methods, including LR and SVM, with L1 and L2 regularization. Table 1 illustrates the model performance across different approaches in the marijuana-323 dataset. Our experiments reveal that among the linear classification methods—LR with L1 regularization, LR with L2 regularization, and SVMs with both L1 and L2—the SVM with L1 regularization achieved the highest classification accuracy at 79.2%. However, the same linear methods did not yield superior results in the HCP dataset. Only SVM with L2 regularization achieved a classification accuracy of 69.6%. These classification results across the two datasets indicate the presence of nonlinear structures among brain regions in the fMRI scans of long-term cannabis users, which linear classifiers fail to discern based on functional connectivity features.
3.2.2 Graph-level classification
We compared the proposed method with other graph-based methods such as GCN, GAT, GraphSAGE, and BrainGB as the baseline models. Table 1 shows that GAT and BrainGB achieve a better area under the curve (AUC) scores than linear classifier methods in the marijuana-323 dataset. However, linear models perform better in accuracy, precision, and recall. However, only GAT performs robust classification in the HCP datasets.
In our ablation experiments, we verified the contribution of the sub-models of the MixHop and LSTM sub-models in our overall model framework. First, we removed the high-order fusion of the MixHop model in the HOGANN and observed a substantial improvement in AUC, reaching 81.9% and 70.4% compared to the graph-based baseline. This indicates that the overall predictive capability of the model is limited when relying solely on the sequential LSTM. Additionally, it compared the BrainGB framework, specifically designed for the brain network, with different message-passing or pooling strategies. The experimental results from BrainGB indicate that utilizing only the one-hop neighborhood of local nodes achieves an AUC of 83.4% in the Marijuana-323 dataset but fails to perform adequately in the HCP dataset. Meanwhile, precision and recall are lower than the original GAT model, with values of 73.6% and 69.3%, respectively. Next, we eliminated the influence of LSTM in LOGAB. It was observed that incorporating MixHop into the model substantially enhanced the AUC for overall brain network learning, achieving 85.1% for marijuana-323 and 75.5% for the HCP dataset, with accuracies of 80.7% and 71.2%, respectively.
One of the advantages of HOGANN is its ability to integrate two sub-models: one dedicated to extracting temporal graph features and another focused on learning higher-order attention, ultimately combining these features for more effective messaging. Experimental results demonstrate its superiority over linear models and other graph-based methods, resulting in increased classification accuracy at 80.7%, precision at 81.0%, and recall at 80.7% compared to state-of-the-art models.
Next, we assessed the cross-validation performance of HOGANN by partitioning the complete dataset into five folds and conducting a 5-fold cross-validation. As depicted in Fig. 7, our HOGANN exhibits a notable AUC of 82.4% $\pm$ 0.029 in multigraph classification.
3.3 Classification based on varying length of fMRI time series
This section explores the impact of segmenting full-time series and subsequences into multigraphs with different window sizes $T^{\prime}$. Our training strategy employs four-time windows of length $T^{\prime}\in$0, 100, 300, 600} for training and testing. As depicted in Fig. 4, it is evident that the AUC, accuracy, F1, and precision exhibit substantial reductions as the number of segmented fMRI signals increases (segment $>$ 3) in the marijuana-323. However, GAT demonstrates a more stable performance when segmenting two or three times compared to GCN. HOGANN showcases increasingly robust classification performance in our experiments as the number of segments increases. Specifically, under the optimal condition of segmenting fMRI signals twice, HOGANN achieves an AUC of 82.7%, accuracy of 79.2%, F1 of 79.26%, and precision of 79.3%.
3.4 Ablation study
We conducted ablation studies on the sub-model effect to explore the impact of the loss function. Several optimization experiments were designed to compare the effectiveness of Eq. (19) on the overall model learning, as presented in Table 2. Subsequently, attention mechanism ablation experiments were conducted on the effect of multi-head attention, with each head independently attending to information from neighboring nodes’ latent space. Following that, we compared the classification effect of different $K$ nearest neighbor parameters on multigraph classification in two datasets.
  1. 1.
    We compared the performance solely using $\mathcal{L}_{c}$. from the experimental results. Our findings demonstrate that HOGANN, which integrates two sub-models, achieved an AUC of 78.8% and 66.1% in two datasets, respectively. Furthermore, using only $\mathcal{L}_{MixGAT}$ for the high-order attention effect showed no difference from $\mathcal{L}_{c}$, and its accuracy, precision, and recall were lower than when using only $\mathcal{L}_{c}$ in marijuana-323. It is evident that $\mathcal{L}_{c}$ can still preserve the optimization effects of high- order attention within a single loss function.
  2. 2.
    We believe that combining and optimizing loss functions may be better than using a single loss function. However, our experiment using a weight of $\alpha=0.3$ did not improve the results. Rather, it simply added the the two loss functions $\mathcal{L}_{c}+\mathcal{L}_{MixGAT}$. This indicates that HOGANN indeed improves optimization by combining the two losses, resulting in a classification AUC of 84.0%, an accuracy of 78.4%, a precision of 78.3%, and a recall of 78.4% in marijuana-323. Similar improvements in performance were seen in HCP.
  3. 3.
    In Figure 8, we increase the number of graph attention multi-heads to capture long-range dependencies and local neighborhood graph structure. From the results, it can be observed that HOGANN in marijuana-323 and HCP substantially improved classification performance. This indicates that increasing the number of attention heads can enable a broader learning of node embeddings for the high-order connectivity of each vertex.
  4. 4.
    We analyzed the HCP dataset characterized by marijuana dependence (Mj) and those who have previously used marijuana, conducting predictions across different age groups. The experimental results indicate that for the group with marijuana dependence, our HOGANN predicts the results more accurately in all age groups compared to the group that has only used marijuana. This indicates that individuals with marijuana dependence exhibit stronger brain region activity in the BOLD fMRI time series.
Table 2: Classification Performance by Age Group in HCP Dataset
Condition Age Group AUC (%) Accuracy (%)
Marijuana Dependence 22-25 88.5 92.4
25-30 89.5 86.4
30-37 98.8 87.9
Ever Used Marijuana 22-25 72.1 81.1
25-30 72.2 63.1
30-37 70.9 65.9
Figure 4: Comparing different models based on their classification performance on the marijuana-323 dataset with varying multi-segment time window sizes.
Figure 5: The average of predicted functional weighted connectivity (WC) and degree centrality (DC) values ranking between ROIs in marijuana-323. Panel (A) illustrates the comparison between the HC and LM groups, while Panel (B) displays the corresponding network of the HC and LM group, highlighting the regions with the highest DC values.
Figure 6: A comparison of machine learning methods with our HOGANN in identifying the optimal community clustering. In Panel (A), we compare community detection in weighted functional connectivity matrices using machine learning and our HOGANN, where we map the most significant four communities of the LM group to their corresponding brain regions of the connectivity network. Based on Panel (B) for the LM group, we depict four communities corresponding to the colored boxes, illustrating the connectivity of the most active brain regions in Panel (C).
Figure 7: Evaluating AUC performance using five-fold cross-validation on the marijuana-323 dataset for multigraph classification.
Figure 8: Panel (A) demonstrates the performance changes with multiple attention heads, while Panel (B) illustrates the classification performance changes when the neighborhood K in of the HOGANN is varied.
Table 3: Loss comparison
Marijuana-323 HCP
Strategy AUC (%) Acc. (%) Prec. (%) Rec. (%) AUC (%) Acc. (%) Prec. (%) Rec. (%)
$\mathcal{L}_{c}$ 78.8 76.9 76.7 76.9 66.1 63.9 63.9 63.9
$\mathcal{L}_{MG}$ 78.9 73.9 73.7 73.9 62.4 68.0 68.3 68.0
$\alpha\mathcal{L}_{MG}+(1-\alpha)\mathcal{L}_{c}$ 78.9 75.4 75.4 75.4 59.2 62.5 62.5 62.5
$\mathcal{L}_{c}+\mathcal{L}_{MG}$ 84.0 78.5 78.3 78.5 70.0 69.1 69.4 69.1
Table 4: Important Brain Connections in Marijuana-323
ROI index Connection Description
(1, 89) “L anterior insula” and “R crus cerebellum - visiospatial”.
(4, 89) “R anterior insula” and “R crus cerebellum - visiospatial”.
(6, 89) “R crus cerebellum - ant_sal” and “R crus cerebellum - visiospatial”.
(7, 89) “L superior temporal/auditory” and “R crus cerebellum - visiospatial”.
(10, 89) “R striatum/thalamus” and “R crus cerebellum - visiospatial”.
(12, 89) “L inferior frontal gyrus” and “R crus cerebellum - visiospatial”.
(16, 27) “L lateral angular gyrus” and “L mid-temporal cortex”.
(22, 89) “L parahippocampal gyrus” and “R crus cerebellum - visiospatial”.
(24, 89) “L mid occipital cortex” and “R crus cerebellum - visiospatial”.
(26, 89) “L inferior frontal gyrus” and “R crus cerebellum - visiospatial”.
(28, 89) “L mid-posterior temporal cortex” and “R crus cerebellum - visiospatial”.
(29, 89) “L medial angular gyrus” and “R crus cerebellum - visiospatial”.
(30, 89) “R inferior frontal gyrus” and “R crus cerebellum - visiospatial”.
(31, 89) “R mid-temporal cortex” and “R crus cerebellum - visiospatial”.
(52, 89) “bilateral medial posterior precuneus” and “R crus cerebellum - visiospatial”.
(55, 89) “bilateral calcarine cortex” and “R crus cerebellum - visiospatial”.
(56, 89) “L LGN” and “R crus cerebellum - visiospatial”.
(63, 64) “L pre/post-central gyri” and “R pre/post-central gyri”.
(67, 89) “cerebellar vermis” and “R crus cerebellum - visiospatial”.
(81, 89) “L precentral/fronto-opercular region” and “R crus cerebellum - visiospatial”.
Figure 9: The top three LM groups activate brain regions with mean DC values across four communities in the Marijuana-323.
Figure 10: HOGANN predicted craving location in the brain (green circle) using carving and gray regions of craving maps, Achieving high accuracy in determining overlaps related to chronic marijuana use in marijuana-323.
Figure 11: The HOGANN predicts a seven-subnetwork parcellation of Yeo in the Marijuana-323.
3.5 Community clustering
This study explores the community structure within brain networks, emphasizing the significance of clustering for each region based on reconstructed brain networks from Eq. (14). Our analysis examines the intricate connections between LM users and HC in the brain. This examination is facilitated by integrating the functional connectivity of each instance with the trainable weighted matrix derived from our HOGANN. Furthermore, we can rank the modularity and degree centrality (DC) by utilizing the average trainable weighted connectivity matrix to identify distinct clusters and community regions and compare with ML and our proposed model as demonstrated in Fig. 5. Further, we can find that the LM groups in Fig. 5 (A) have more significant community clustering, and we find out the significant region of the top five DC values in Fig. 5 (B).
We conducted a comparative analysis between traditional linear machine learning methods and our proposed HOGANN. This comparison, illustrated in Fig. 6 (linear methods) and B (HOGANN), highlights the superiority of the HOGANN in achieving enhanced modularity and clustering effects, particularly in community detection, as substantiated by the results.
As depicted in Fig. 6 (C), we encircled squares based on the color of each brain cluster region, representing their high-density degree of centrality. By identifying critical brain regions associated with four communities and sorting the importance scores according to the average degree centrality values, we highlighted them. This allowed us to pinpoint essential brain regions linked to four distinct communities. In Table 4, we have followed the indices corresponding to 90 ROIs from Stanford Atlas definition  [36] and calculate DC scores that can be found within the four communities, as shown in Fig. 9. It reveals the importance of the rankings of these values, offering insights into the structural and functional connectivity across these brain regions. In our section 3.7 on brain analysis, we will analyze and interpret the subnetworks corresponding to the four communities through Yeo’s seven subnetworks.
3.6 Craving maps characterization
Craving maps can aid in deciphering brain network alterations in long-term marijuana users. We utilized the Neurosynth package (https://neurosynth.org/), aligning it with the ”craving” keyword to identify consistently active brain regions across studies strongly associated with the term “craving” [41]. Leveraging the design of the cited study, an automated meta-analysis of 80 publications identified regions activated by ”craving,” and Stanford ROIs can be mapped to the uniformity map  [33]. Specifically, the uniformity map measures low-activity voxels by applying a threshold of 25% during processing and identifies the activity in each brain region. The identified ROIs on the resulting map can be marked in red as threshold activity areas, as shown in Fig. 10. These ROIs represent areas substantially associated with craving, offering a comprehensive view of the brain’s response to LM use.
Next, we utilized the HOGANN to predict craving regions based on degree centrality (depicted by the green circle). We use uniformity maps to compare these predicted regions to the Neurosynth meta-analytic database. Finally, we invert the y-axis on the brain volumes within the Stanford ROIs, as indicated by the gray circle. The outcomes shown in Fig. 10 demonstrate the concordance of our HOGANN in predicting green cravings that are closely aligned with the ROIs across the axes at y = -60, y = 50, y = 32, y = 44 and y = 58. Our results showcase a successful analysis of cognitive-behavioral regions in LM users by effectively predicting specific activity regions. The prediction results through our model improve the prediction consistency and concordance of predictions between the network region and the craving maps from the meta-analysis of existing studies.
3.7 Alterations in the intrinsic functional connectivity
Yeo et al. [42] identified seven intrinsic functional brain networks of the cerebral cortex using the cluster approach based on resting state fMRI data from 1000 participants [42]. We utilized Yeo’s atlas of seven brain networks and analyzed the signal changes among these networks for SUD patients and normal controls. The seven networks include: 1. Visual attention 2. Somatosensory 3. Dorsal attention 4. Ventral attention 5. Limbic 6. Frontoparietal 7. Default mode network. In Fig. 11, we plotted a uniformity map based on Yeo 7 resting-state subnetworks and marked regions predicted by HOGANN (in blue). In the first row, the green regions correspond to overlapping subnetworks at the top of the atlas, which are areas of the Frontoparietal network associated with cognitive functions such as decision-making, problem-solving, planning, and executive control. In the second row, purple regions represent dorsal attention (attention and cognitive control). With Yeo’s seven subnetworks, we found the HOGANN could identify long-term marijuana users with substantial regions in an active transition of cognitive function and attention.
By comparing Yeo’s seven subnetworks to the four communities in Fig. 9, we can see that mean DC well explains the top three regions of each community. Communities 1 and 2 correspond to the dorsal attention network. Communities 4 and 5 correspond to the frontoparietal control network. Finally, Community 3 corresponds to regions that are either part of or closely related to the default mode network. According to our study’s observations, we align with previous relevant research that demonstrates LM users exhibit increased subcortical hyperconnectivity, leading to a widespread increase in functional connectivity across the whole brain connectome, particularly in the dorsal attention network and the default mode network [15, 43].
4 Conclusion
This study proposes a mixed model (HOGANN), which incorporates a dynamic graph LSTM capable of learning temporal patterns in fMRI time series and integrates a high-order attention mechanism. Moreover, the HOGANN is more accurate than other models for predicting marijuana addiction based on the multigraph of brain networks. By utilizing the HOGANN and Yeo atlas of seven intrinsic networks, we identified the most relevant subnetworks and cognitive regions negatively influenced by persistent marijuana use, revealing that chronic marijuana use adversely affects cognitive control, particularly within the dorsal attention and frontoparietal networks, which are essential for attentional, cognitive control, and higher cognitive functions. Therefore, through the HOGANN, we can more accurately predict craving maps in LM users, identify brain maps associated with long-term cravings, and pinpoint active areas important for analysis. In the future, our study not only deepens our comprehension of the effects of long-term marijuana use on the brain but also underscores the revolutionary potential of AI in analyzing functional brain networks.